OpenStudy (anonymous):

find the missing number 1,12,54,160,375,?,1372 it involves both square numbers and triangular numbers

OpenStudy (campbell_st):

well the triangular numbers are 1, 3, 6, 10, 15, 21, 28... the square numbers are 1, 4, 9, 16, 25, 36, 49, ... the numbers in the sequence are triangular numbers multiplied by the equivalent square number \[1 \times 1, 3 \times 4, 6 \times 9, 10 \times 16..... \] hope it helps

OpenStudy (campbell_st):

so I thin you need the 6th triangular number multiplied by the 6th square number